### v3

# Measuring Social Biases in Masked Language Models by Proxy of Prediction Quality

Abstract
Transformer language models have achieved state-of-the-art performance for a variety of natural language tasks but have been shown to encode unwanted biases. We evaluate the social biases encoded by transformers trained with the masked language modeling objective using proposed proxy functions within an iterative masking experiment to measure the quality of transformer models’ predictions and assess the preference of MLMs towards disadvantaged and advantaged groups. We find all models encode concerning social biases. We compare bias estimations with those produced by other evaluation methods using benchmark datasets and assess their alignment with human annotated biases. We extend previous work by evaluating social biases introduced after retraining an MLM under the masked language modeling objective and find proposed measures produce more accurate and sensitive estimations of biases introduced by retraining MLMs based on relative preference for biased sentences between models, while other methods tend to underestimate biases after retraining on sentences biased towards disadvantaged groups.
Warning: This paper contains explicit statements of biased stereotypes and may be upsetting.
1 Introduction
Masked language models (MLM(s); Brown et al., 2020; Devlin et al., 2019; Liu et al., 2019a; Radford et al., 2019) such as transformers BERT and RoBERTa incorporate bidirectionality and self-attention, producing contextually-aware embeddings. Unlike $\text{BERT}_{\text{unc}}$, RoBERTa was pretrained solely on the masked language modeling objective (MLMO) on a larger corpus of text. RoBERTa uses a dynamic masking strategy for a diverse set of representations during training and has achieved state-of-the-art results on GLUE, RACE, and SQuAD (Liu et al., 2019b; Rajpurkar et al., 2016; Lai et al., 2017; Wang et al., 2018). Distilled model variants have been shown to train significantly faster for minor decreases in performance Sanh et al. (2019). MLMs have produced state-of-the-art results for masked language modeling, named entity recognition, and intent or topic classification, but also encode concerning social biases against disadvantaged groups that are undesirable in production settings. As MLMs become increasingly prevalent, researchers have been working on methods to measure biases embedded in these models Nangia et al. (2020); Kaneko and Bollegala (2022); Salutari et al. (2023).
To address the issue of social biases in MLMs at its source, we measure biases of MLMs while focusing on an MLM’s key pretraining objective: masked language modeling. In this work, we focus on proposing and assessing bias evaluation measures for MLMs and not proposing methods for debiasing MLMs. However, when assessing evaluation measures, we consider prior research that involves retraining or fine-tuning under the MLMO with debiased or counterfactual data to reduce biases in MLMs, such as Zhao et al., 2018 and Zhao et al., 2019 which use data augmentation to swap gendered words with their opposites prior to retraining. Motivated by this, we assess whether proposed measures satisfy an important criterion for improvement over previously proposed ones: alignment with biases introduced by MLM retraining under the MLMO (Section 4.5.1).
We represent MLM bias through a model’s relative preference for ground truth tokens between two paired biased sentences, where one contains bias against disadvantaged groups (stereotypical) and the other contains bias against advantaged groups (anti-stereotypical). We assess the relative preferences of MLMs towards sentences biased against disadvantaged and advantaged groups using proxy measures for prediction quality. In particular, we quantify MLM preference by proxy of masked token prediction quality given unmasked token context between encoded sentences within pairs.
To measure the preference of an MLM using the (attention-weighted) quality of its predictions under the MLMO, we propose and validate proxy functions $\Delta\textsc{pa}$ (Equation 7; Probability Difference with Attention) and crra (Equation 6; Complementary Reciprocal Rank with Attention) and introduce a modified $\Delta\textsc{p}$ (Equation 3) to measure the likelihood an MLM will select a ground truth token to replace a masked one, and extend these definitions for a sentence. We apply per-model indicator function bspt (Equation 9) to estimate the encoded social biases in pretrained MLMs. Our approach differs from prior research in measuring biases in MLMs by using attention-weights under an Iterative Masking Experiment (ime; Section 4) to probe MLM preferences.
We compare pre- and retrained MLMs within the same model class to recover the nature of biases introduced by retraining. In particular, we define and apply a proxy for the relative preference between two MLMs with model-comparative indicator function bsrt (Bias Score for MLM Re-training; Equation 10) to estimate social biases introduced by retraining MLMs under the MLMO. [^1] We acknowledge that ”introduced” biases must be represented by the relative change in biases after retraining an MLM under the MLMO.
In summary, the primary contributions of this work are as follows:
  • We explore MLM bias through a model’s relative preference for ground truth tokens between two minimally distant sentences with contrasting social bias under the ime. We measure this using the attention-weighted quality of predictions.
  • We propose model-comparative indicator function bsrt to estimate social biases against disadvantaged groups for a retrained MLM relative to its pretrained base, and assess bias evaluation measures for alignment with biases introduced by MLM retraining under the MLMO.
  • We evaluate social biases for four transformer models available through the Transformers library Wolf et al. (2020). We use proxy measures for MLM prediction quality with model-comparative function bsrt to estimate social biases introduced by retraining MLMs under the MLMO. We find proposed measures $\Delta\textsc{pa}$ and crra, along with modified $\Delta\textsc{p}$ and crr, produce more accurate estimations of biases introduced by MLM retraining than previously proposed ones, which can produce concerning underestimations of biases after retraining MLMs on sentences biased against disadvantaged groups. We observe that proposed measures $\Delta\textsc{pa}$ and crra, along with modified $\Delta\textsc{p}$, show greater sensitivity than crr, csps, aul, and aula to relative changes in MLM bias due to retraining, indicated by larger and smaller bias scores and more frequently significant relative difference in proportions of bias between pre- and retrained transformers.
Our methodology could help others evaluate social biases encoded in an MLM after it is retrained on the MLMO, such as for any downstream fill-mask task, or after it is fine-tuned for other objectives that alter weights and change MLMO performance. We release a package for measuring biases in MLMs to enable bias score computation on user-supplied or benchmark datasets and easy integration into existing evaluation pipelines.
2 Related Work
2.1 Biases in Static Word Embeddings
Static word embeddings Pennington et al. (2014); Mikolov et al. (2013) can be shifted in a direction to decompose bias embedded in learned text data representations. These could be analogies or biases along an axis, such as gender (Bolukbasi et al., 2016) or race (Manzini et al., 2019). The Word Embedding Association Test (WEAT; Caliskan et al., 2017) measures association between targets and attributes using cosine similarity between static word embeddings, but has been shown to overestimate biases by Ethayarajh et al., 2019 which proposed the robust relational inner product association (RIPA) method, derived from the subspace projection method to debias vectors in Bolukbasi et al., 2016.
While WEAT has shown that token-level embeddings produced by GloVe and Word2Vec encode biases based on gender and race Caliskan et al. (2017), the Sentence Encoder Association Test (SEAT; May et al., 2019) extends on WEAT to measure social biases in sentence-level encoders such as ELMo and BERT using template sentences with masked target tokens, averaging token embeddings to form sentence-level embeddings on which cosine similarity is applied as a proxy for semantic association. As an alternative evaluation method under a different objective, Liang et al., 2020 assessed differences in log-likelihood between gender pronouns in a template sentence where occupations can uncover the directionality of the bias encoded by an MLM.
2.2 Evaluating Biases in MLMs
When considering a sentence $s$ containing tokens $\{t_{1},t_{2},...,t_{l_{s}}\}$, where $l_{s}$ is the number of tokens in $s$, (modified) token(s) of $s$ can characterize its bias towards either disadvantaged or advantaged groups. For a given sentence $s$ with tokens $t\in s$ we denote all tokens besides $t_{x}$ as $s_{\setminus t_{x}}$ (where $1\leq x\leq l_{s}$), and we denote modified tokens as $M$ and unmodified tokens as $U$ ($s=U\cup M$). For a given MLM with parameters $\theta$, we denote a masked token as $t_{m}$ and a predicted token as $t_{p}$.
Salazar et al., 2020 uses pseudo-log-likelihood scoring to approximate $P(U|M,\theta)$, or the probability of unmodified tokens conditioned on modified ones. Similarly, Nangia et al., 2020 reports CrowS-Pairs Scores (csps; Appendix B), a pseudo-log-likelihood score for an MLM selecting unmodified tokens given modified ones. Nadeem et al., 2021 reports a StereoSet Score (sss; Appendix C), a pseudo-log-likelihood score for an MLM selecting modified tokens given unmodified ones.
Salutari et al., 2023 tests MLMs in an iterative fill mask setting where the model outputs a set of tokens (or the (log)softmax of model logits mapped to tokens) to fill the masked one, starting with the token of highest probability $P(t_{p}|c)$ and, as such, first rank $\rho(t_{p}|c)=1$ in the set of model token predictions (limited by the MLM’s embedding space).
$\begin{split}\Delta\textsc{p}(t|s_{\setminus t_{m}};\theta)&=P(t_{p}|s_{\setminus t_{m}};\theta)-P(t_{m}|s_{\setminus t_{m}};\theta)\\ &=\Delta\textsc{p}(w;\theta)\end{split}$ (1)
$\Delta\textsc{p}(t|s_{\setminus t_{m}};\theta)$ (Equation 1) represents the difference between the probability of a predicted token $t_{p}$ and a masked token $t_{m}$ in a sentence $s$. It serves as a proxy of the MLM’s prediction quality for a token given its context within the ime, or all tokens in $s$ besides $t_{m}$ (Salutari et al., 2023).
$\begin{split}\textsc{crr}(t|s_{\setminus t_{m}};\theta)&=\bigl{(}\rho(t_{p}|s_{\setminus t_{m}};\theta)^{-1}-\rho(t_{m}|s_{\setminus t_{m}};\theta)^{-1}\bigr{)}\\ &=1-\rho(t_{m}|s_{\setminus t_{m}};\theta)^{-1}\\ &=\textsc{crr}(w;\theta)\end{split}$ (2)
Salutari et al., 2023 proposes the Complementary Reciprocal Rank ($\textsc{crr}(t|s_{\setminus t_{m}};\theta)$; Equation 2) for a masked token given its context, where $\rho(t_{p}|s_{\setminus t_{m}})^{-1}$ is the reciprocal rank of the predicted token (and always equal to 1) and $\rho(t_{m}|s_{\setminus t_{m}})^{-1}$ is the reciprocal rank of the masked token. Thus, $\rho(t_{m}|s_{\setminus t_{m}})^{-1}$ provides a likelihood measure for $t_{m}$ being chosen by the model as a candidate token to replace the ground truth (masked) one.
Salutari et al., 2023 defines $\Delta\textsc{p}(s)$ (Appendix D) as the probability difference for a sentence $s$ and $\textsc{crr}(s)$ (Appendix E; average of all single masked token crrs with token ordering preserved) as the complementary reciprocal rank for a sentence $s$, and uses them as proxies for an MLM’s prediction quality or preference.
Kaneko and Bollegala, 2022 propose evaluation metrics All Unmasked Likelihood (aul; Appendix F) and aul with Attention weights (aula; Appendix G), where both are generated by requiring the MLM to predict all tokens (unmasked input). By requiring the MLM to simultaneously predict all of the tokens in a given unmasked input sentence $s$, Kaneko and Bollegala, 2022 aim to eliminate biases associated with masked tokens under previously proposed pseudo-likelihood-based scoring methods (Nadeem et al., 2021, Nangia et al., 2020), which assumed that masked tokens are statistically independent, and selectional biases from masking a subset of input tokens, such as high frequency words (masked more often during training).
Kaneko and Bollegala, 2022 reference the use of sentence-level embeddings produced by MLMs for downstream tasks such as sentiment classification to argue that biases associated with masked tokens should not influence the intrinsic bias evaluation of an MLM, as opposed to the evaluation of biases introduced after an MLM is fine-tuned. In contrast, we focus on an MLM’s key pretraining objective, masked language modeling, to measure social biases of the MLM. [^2] Masked language modeling was a pretraining objective for all transformers considered in this work. Next sentence prediction was not used for RoBERTa and variants due to relatively lower performance with its inclusion Liu et al. (2019b). In addition, we measure relative changes in biases w.r.t. the intrinsic biases of the same base MLM after retraining under the MLMO (as opposed to fine-tuning). Thus, we argue that biases associated with masked tokens are not undesirable in our case.
aul and aula were found to be sensitive to contextually meaningful inputs by randomly shuffling tokens in input sentences and comparing accuracies with and without shuffle. crr is also conditional to the unmasked token context by definition. We argue measure sensitivity to unmasked token contexts is desirable when evaluating a given MLM’s preference under a fill-mask task, or when estimating biases using contextualized token-level embeddings produced by an MLM.
In contrast with previous methods such as sss and csps, we refrain from using strictly modified or unmodified subsets of input tokens as context and instead provide all tokens but the ground truth one as context for MLM prediction under each iteration of our masking experiment. In this sense, crr and $\Delta\textsc{p}$ consider all tokens equally, and like aul, they might also benefit from considering the weight of MLM attention as a proxy for token importance when probing for MLM preferences for ground truth tokens between two paired sentences along a social bias axis. [^3] By incorporating attention weights, $\Delta\textsc{pa}$ and crra consider token-level contributions to MLM predictions, leveraging the attention mechanism’s role in weighting contextually important tokens.
Existing benchmark datasets such as CPS are limited to one ground truth per masked token, so an important consideration is an MLM’s ability to predict multiple plausible tokens for a context that could qualify for concerning social bias but goes unrecognized during evaluation using previously proposed measures. crr could perform better than pseudo-(log)likelihood-based measures for this sensitivity that yields larger relative differences in $\Delta\text{P}(t|c)$ as opposed to $\textsc{crr}(t|c)$ [^4] The lowest rank of any possible ground truth token within a model’s vocabulary is equal to the size of the vocabulary ($1\leq\rho(t_{m}|c)\leq V$ and $0\leq\textsc{crr}(s)\leq 1-(V)^{-1}$ for an MLM with embedding vocabulary of size $V$)., and is deemed critical for evaluation by Salutari et al., 2023 due to the possible uniformity of probabilities generated by a particular MLM w.r.t. others.
3 Methodology
3.1 Biases in Pretrained MLMs
Model ime Input with masked token $w_{11}$ = money $w_{11}$ Rank
$\text{distilRoBERTa}_{P}$ The poor are really ignorant about how to handle the problems they do have 5
$\text{distilRoBERTa}_{R}$ The poor are really ignorant about how to handle the money they do have 1
$\text{distilRoBERTa}_{P}$ The rich are really ignorant about how to handle the wealth they do have 4
$\text{distilRoBERTa}_{R}$ The rich are really ignorant about how to handle the problems they do have 2
Table 1: MLM predictions and ground truth masked token ranks for input sentences under the ime. The processed input sentences are "The [poor, rich] are really ignorant about how to handle the money they do have", and are contained in $S_{\text{dis}}$ or $S_{\text{adv}}$ from CPS with characteristic words poor and rich respectively. Masked tokens are in bold.
We probe for MLM preferences using the ime, which masks one token at a time until all tokens have been masked, or until we have $n$ logits or predictions for a sentence with $n$ tokens. Appendix I shows an ime example for one model and text.
Table 1 shows masked token predictions (those with first rank and highest probability) produced by MLMs for example input contexts. The ground truth masked token rank, or $\rho(t_{m}|c)$, represents the quality of the models’ predictions relative to a sentence biased against an advantaged or disadvantaged group. Special start and end character tokens for MLMs are not included in the span of tokens considered in our experiments to eliminate noise. [^5] Special start and end character tokens for MLMs are not considered by measures using attention weights and probabilities computed from the (log)softmax of model logits for the masked token index.
$\begin{split}\Delta\text{P}(t|s_{\setminus t_{m}};\theta)&=\bigl{(}\log P(t_{p}|s_{\setminus t_{m}};\theta)-\log P(t_{m}|s_{\setminus t_{m}};\theta)\bigr{)}\\ &=\Delta\textsc{P}(w;\theta)\end{split}$ (3)
Motivated by the typically long-tail distribution of probabilities output by MLMs, we redefine Equation 1 and propose a modified version $\Delta\text{P}(w)$ as shown in Equation 3. The log transformation reduces sensitivity to differences when probabilities are high but enhances relative differences when probabilities are low, improving sensitivity to subtle but potentially meaningful variations in low-probability token choices.
$\begin{split}\textsc{crra}(t|s_{\setminus t_{m}};\theta)&=a_{m}\bigl{(}1-\log\rho(t_{m}|s_{\setminus t_{m}};\theta)^{-1}\bigr{)}\\ &=\textsc{crra}(w;\theta)\end{split}$ (4)
$\begin{split}\Delta\textsc{pa}(t|s_{\setminus t_{m}};\theta)&=a_{m}\bigl{(}\log P(t_{p}|s_{\setminus t_{m}};\theta)\\ &\hskip 32.0pt-\log P(t_{m}|s_{\setminus t_{m}};\theta)\bigr{)}\\ &=\Delta\textsc{pa}(w;\theta)\end{split}$ (5)
For a given MLM with parameters $\theta$, we propose attention-weighted measures $\textsc{crra}(w)$ and $\Delta\text{PA}(w)$ defined in Equations 4 and 5 respectively, where $a_{w}$ is the average of all multi-head attentions associated with the ground truth token $w$ [^6] Averaging the multi-head attentions associated with a ground truth token provides a potentially balanced representation of token-level contributions by capturing diverse contextual relationships encoded by individual heads while reducing noise from any single head., and $P(t_{m}|s_{\setminus t_{m}})$ and $\rho(t_{m}|s_{\setminus t_{m}})$ are the probability score and rank of the masked token respectively. We extend these definitions for a sentence $s$ as shown in Equations 6 and 7.
$\textsc{crra}(s):=\frac{1}{l_{s}}\sum_{w\in s}\textsc{crra}(w;\theta)$ (6)
$\Delta\textsc{pa}(s):=\frac{1}{l_{s}}\sum_{w\in s}\Delta\textsc{pa}(w;\theta)$ (7)
[…]
$\Delta f_{T}(i)=\begin{cases}f_{T}(S_{\text{adv}}(i))-f_{T}(S_{\text{dis}}(i)),&\text{if }f\in M_{1}\\ f_{T}(S_{\text{dis}}(i))-f_{T}(S_{\text{adv}}(i)),&\text{if }f\in M_{2}\end{cases}$ (8)
[…]
$\textsc{bspt}(f):=\frac{100}{N}\sum_{i=1}^{N}\mathds{1}\bigl{(}\Delta f_{T}(i)>0\bigr{)}$ (9)
$\mathds{1}$ is a per-model indicator function which returns $1$ if transformer $T$ has a larger preference for $S_{\text{dis}}$ relative to $S_{\text{adv}}$ and 00 otherwise, as estimated for a measure $f$ by Equation 8. bspt represents the proportion of sentences with higher relative bias against disadvantaged groups for a given measure, where values above 50 indicate greater relative bias against disadvantaged groups for an MLM.
3.2 Biases Introduced by MLM Retraining
We define a bias score for a retrained MLM (relative to its pretrained base) as bsrt (Equation 10).
$\textsc{bsrt}(f):=\frac{100}{N}\sum_{i=1}^{N}\mathds{1}\bigl{(}\Delta f_{T_{1}}(i)>\Delta f_{T_{2}}(i)\bigr{)}$ (10)
We define a proxy for the relative preference between two MLMs with model-comparative indicator function $\mathds{1}$, which returns $1$ if transformer $T_{1}$ has a larger preference for $S_{\text{dis}}$ relative to $S_{\text{adv}}$ than transformer $T_{2}$ and 00 otherwise, as estimated for a measure $f$ by Equation 8. bsrt can be applied to compare pre- and retrained MLMs within the same model class and recover biases introduced by MLM retraining. Values above 50 indicate greater bias against disadvantaged groups for transformer $T_{1}$ relative to $T_{2}$.
4 Experiments and Findings
4.1 Benchmark Datasets for Social Bias
The Crowdsourced Stereotype Pairs Benchmark (CPS; Nangia et al., 2020) contains biased sentences towards historically advantaged and disadvantaged groups along nine forms of social biases. StereoSet (SS; Nadeem et al., 2021) contains intrasentence and intersentence (with context) pairs for four forms of social biases, using the likelihood of modified tokens given unmodified token contexts as proxy for MLM preference. Similarly, CPS contains characteristic words that distinguish sentences within pairs and define the nature of a particular bias towards either advantaged or disadvantaged groups, but instead uses the relative likelihood of unmodified tokens being chosen by the MLM given a modified context (characteristic word) across a sentence pair (see Appendix A for details on bias categories and sentence counts).
To probe for biases of interest that are encoded in MLMs, the scope of our experiments include all bias categories and sentences pairs in CPS and intrasentence pairs in SS, since intersentence pairs are not masked for bias evaluation (Kaneko and Bollegala, 2022). We estimate MLM preference towards a stereotypical sentence over a less stereotypical one for each bias category in CPS and SS.
4.2 Retraining Dataset
CPS provides a more diverse alternative to biases expressed by sentence pairs in SS. Biases widely acknowledged in the United States are well represented in CPS [^9] CPS categories are a ”narrowed” version of the US Equal Employment Opportunities Commission’s list of protected categories Nangia et al. (2020). compared to SS, and there is greater diversity of sentence structures in CPS Nangia et al. (2020). CPS has been found to be a more reliable benchmark for pretrained MLM bias measurement than SS, and the validation rate of CPS is 18% higher than SS. Nangia et al. (2020). Based on these findings, paired with (1) the computational expense and time-consumption involved with retraining MLMs under the MLMO and (2) concerns regarding standard masked language modeling metric viability on SS, we proceed to use sentence sets in CPS to re-train MLMs and validate methods for estimating the biases that are introduced Nangia et al. (2020) (see Appendix A.2 for details). We re-train MLMs $\forall s\in S_{\text{dis}}$ or $\forall s\in S_{\text{adv}}$, where $s$ is a sentence in CPS biased towards either advantaged ($S_{\text{adv}}$) or disadvantaged groups ($S_{\text{dis}}$).
4.3 Transformer-based Language Models
We denote pre- and retrained transformers as $T_{P}$ and $T_{R}$ respectively. The subscript ${}_{\text{unc}}$ denotes an uncased model. We report results from the following transformer-based language models available through the HuggingFace library Wolf et al. (2020): bert-base-uncased ($\text{BERT}_{\text{unc}}$; Devlin et al., 2019), roberta-base (RoBERTa; Liu et al., 2019b), distilbert-base-uncased ($\text{distilBERT}_{\text{unc}}$; Sanh et al., 2019), and distilroberta-base (distilRoBERTa; Liu et al., 2019a).
4.4 Pretrained MLM Bias Scores
We use bspt to compare MLM preferences, report overall bias scores in Table 2, and provide detailed results for all measures and MLMs in Appendix K.
[…]
CPS Dataset
csps 59.35 60.48 59.35 56.83
aul 58.75 48.34 53.32 51.59
aula 58.09 48.21 51.86 52.65
crr 58.89 61.07 57.76 56.23
crra 60.68 58.89 61.94 60.08
$\Delta\textsc{p}$ 59.88 60.08 59.75 57.49
$\Delta\textsc{pa}$ 60.15 60.81 59.81 58.02
SS Dataset
sss 61.06 59.16 61.4 60.59
aul 59.45 48.91 60.21 51.71
aula 58.83 50.28 59.59 51.66
crr 57.83 53.85 54.37 53.42
crra 62.06 58.59 60.54 61.11
$\Delta\textsc{p}$ 62.2 58.64 61.4 59.31
$\Delta\textsc{pa}$ 62.35 58.21 61.35 59.31
[…]
Overall, all evaluation methods show concerning social biases against disadvantaged groups embedded in MLMs as observed in prior research (Kaneko and Bollegala, 2022, Nangia et al., 2020). Interestingly, $\text{BERT}_{\text{unc}}$ has the lowest overall sss, aul, aula, crra, $\Delta\textsc{p}$, and $\Delta\textsc{pa}$ (second lowest crr) on SS, but conflicting results on CPS, where it has the highest csps, crr, $\Delta\textsc{p}$, and $\Delta\textsc{pa}$ but the lowest aul, aula, and crra. RoBERTa and distilRoBERTa have higher overall bias than $\text{BERT}_{\text{unc}}$ and $\text{distilBERT}_{\text{unc}}$ according to (1) all but one measure on SS and (2) aul and crra on CPS.
[…]
4.4.1 Human Annotated Bias Alignment
We compare the alignment (agreement) between measures and bias ratings in CPS. Sentence pairs in CPS received five annotations in addition to the implicit annotation from the writer. We map sentences to a binary classification task, where a sentence is considered biased if it satisfies criteria from Nangia et al., 2020, where (1) at least three out of six annotators (including the implicit annotation) agree a given pair is socially biased and (2) the majority of annotators who agree a given pair is socially biased agree on the type of social bias being expressed. [^11] This experiment setting gives 58 unbiased pairs and 1,450 biased pairs for binary classification.
We compute evaluation measures derived from MLMs to predict whether a pair is biased or unbiased at varying thresholds. All measures are computed for each sentence in a pair. Thresholds for bias scores computed on sentences with bias against advantaged and disadvantaged groups respectively maximize area under the ROC Curve for that measure. We find one or more of proposed measures $\Delta\textsc{pa}$ and crra, along with modified $\Delta\textsc{p}$ and crr, outperform aul and aula in their agreement with human annotators on CPS for $\text{RoBERTa}_{P}$ and $\text{BERT}_{P,\text{unc}}$ based on higher AUROC values if MLMs exhibit bias towards disadvantaged groups (ROC curves in Appendix M).
4.5 Retrained MLM Bias Scores
Model Retrain dataset $f:\min/\max\textsc{bsrt}\ \exists c\in C$ $f:p<a\ \forall c\in C$
$\text{RoBERTa}_{R}$ $\forall s\in S_{\text{dis}}$ $\max{\textsc{bsrt}}:\{\textbf{$\Delta\textsc{p}$},\textbf{$\Delta\textsc{pa}$}\}$ crr, crra, $\Delta\textsc{p}$, $\Delta\textsc{pa}$
$\forall s\in S_{\text{adv}}$ $\min{\textsc{bsrt}}:\{\textbf{{crra}},\textbf{$\Delta\textsc{p}$}\}$ csps, crr, crra, $\Delta\textsc{p}$, $\Delta\textsc{pa}$
$\text{distilRoBERTa}_{R}$ $\forall s\in S_{\text{dis}}$ $\max{\textsc{bsrt}}:\{\textbf{{crr}},\textbf{$\Delta\textsc{p}$},\textbf{$\Delta\textsc{pa}$}\}$ $\Delta\textsc{p}$, $\Delta\textsc{pa}$
$\forall s\in S_{\text{adv}}$ $\min{\textsc{bsrt}}:\{\textbf{{crra}},\textbf{$\Delta\textsc{p}$},\textbf{$\Delta\textsc{pa}$}\}$ crr, crra, $\Delta\textsc{p}$, $\Delta\textsc{pa}$
$\text{BERT}_{R}$ $\forall s\in S_{\text{dis}}$ $\max{\textsc{bsrt}}:\{\textbf{$\Delta\textsc{p}$},\textbf{$\Delta\textsc{pa}$}\}$ crr, crra, $\Delta\textsc{p}$, $\Delta\textsc{pa}$
$\forall s\in S_{\text{adv}}$ $\min{\textsc{bsrt}}:\{\textbf{{crra}},\textbf{$\Delta\textsc{p}$},\textbf{$\Delta\textsc{pa}$}\}$ csps, crr, crra, $\Delta\textsc{p}$, $\Delta\textsc{pa}$
$\text{distilBERT}_{R}$ $\forall s\in S_{\text{dis}}$ $\max{\textsc{bsrt}}:\{\textbf{{aul}},\textbf{{crr}},\textbf{$\Delta\textsc{p}$},\textbf{$\Delta\textsc{pa}$}\}$ aul, aula, $\Delta\textsc{p}$, $\Delta\textsc{pa}$
$\forall s\in S_{\text{adv}}$ $\min{\textsc{bsrt}}:\{\textbf{{crra}},\textbf{$\Delta\textsc{p}$},\textbf{$\Delta\textsc{pa}$}\}$ csps, crra, $\Delta\textsc{p}$, $\Delta\textsc{pa}$
Table 3: Measures $\forall f\in\{\textsc{crr},\textsc{crra},\Delta\textsc{p},\Delta\textsc{pa},\textsc{csps},\textsc{aul},\textsc{aula}\}$ with at least one $\min\textsc{bsrt}$ or $\max\textsc{bsrt}$ across bias categories in CPS for MLMs retrained on $S_{\text{adv}}$ or $S_{\text{dis}}$ respectively (bsrt; Equation 10). $f:p<a\ \forall c\in C$ indicates that, for a measure $f$ and every bias category in CPS, the relative difference in proportions of bias between pre- and retrained transformers is statistically significant according to McNemar’s test (p-value < 0.05).
We re-train each transformer under consideration using the PyTorch Python library with P100 and T4 GPUs on cased (RoBERTa and distilRoBERTa) and uncased ($\text{BERT}_{\text{unc}}$ and $\text{distilBERT}_{\text{unc}}$) versions of CPS sentences (see Appendix P.1 for more details).
Figure 1: Difference between ECDFs for crr distribution for sentences in $S_{\text{dis}}$ and $S_{\text{adv}}$ for pre- and retrained RoBERTa and the race bias category in CPS. The line at $y=0$ separates what is biased against disadvantaged groups on the positive y-axis from what is biased against advantaged groups on the negative y-axis.
Figure 1 shows the difference in ECDFs for the $\textsc{crr}_{m}(s)$ measure on the race bias category in CPS, where $m$ is retrained MLM $\text{RoBERTa}_{R}$. It illustrates how the difference in ECDF distributions for an MLM can visually represent a contextual shift in relative bias. When compared with pretrained RoBERTa (RoBERTa-P), we observe an upwards shift in the difference of ECDFs for the crr difference between sentence sets after retraining RoBERTa on $S_{\text{dis}}$ (RoBERTa-R-DIS), and a downwards shift after retraining on $S_{\text{adv}}$ (RoBERTa-R-ADV). This is expected since retraining on $S_{\text{dis}}$ or $S_{\text{adv}}$ should shift the MLM preference towards the corresponding bias type.
We compute bsrt for retrained MLMs and all measures and bias categories on CPS and report detailed results in Appendix P.2.121212We assess whether the relative differences in proportions of bias between pre- and retrained transformers are statistically significant according to McNemar’s test McNemar (1947) (p-value ¡ 0.05), using binarized outcomes for bias as given by $f_{R}(S_{\text{adv}})>f_{R}(S_{\text{dis}})$ and $f_{P}(S_{\text{adv}})>f_{P}(S_{\text{dis}})$ to create contingency tables of outcome pairings between pre- and retrained transformers to test for marginal homogeneity. In general, each measure produces a bias score in accordance with the particular retraining dataset used for almost all significant results across MLMs, demonstrating that proposed function bsrt can be applied to estimate the bias against disadvantaged groups for transformer $T_{1}$ relative to its pretrained base $T_{2}$.
In Table 3, we report (1) measures with at least one $\min\textsc{bsrt}$ or $\max\textsc{bsrt}$ across bias categories in CPS for MLMs retrained on $S_{\text{adv}}$ or $S_{\text{dis}}$ respectively, and (2) measures for which the relative difference in proportions of bias between pre- and retrained transformers is statistically significant according to McNemar’s test (p-value < 0.05) for every bias category in CPS.
For all MLMs retrained on $S_{\text{dis}}$, crr, $\Delta\textsc{p}$, and $\Delta\textsc{pa}$ give results above 50 for each bias category as expected, with higher scores than other measures in almost every case. crra also gives results above 50 for each bias category except for disability bias using $\text{distilBERT}_{R}$. Similarly, for all MLMs retrained on $S_{\text{adv}}$, crr, crra, $\Delta\textsc{p}$, and $\Delta\textsc{pa}$ give results below 50 for each bias category as expected, with lower scores than other measures in almost every case. In contrast, aula and csps give 3 and 4 bias scores below 50 respectively across all MLMs retrained on $S_{\text{dis}}$. Notably, aul and crra give 1 bias score below 50. aul and aula give 4 and 5 bias scores above 50 respectively across all MLMs retrained on $S_{\text{adv}}$.
As shown in Table 3, in every case and across MLMs, one or more of proposed measures reports the highest retraining bias scores for MLMs retrained on $S_{\text{dis}}$ and the lowest for MLMs retrained on $S_{\text{adv}}$. In addition, $\Delta\textsc{p}$ and $\Delta\textsc{pa}$ give significant results for every bias type across all MLMs retrained on $S_{\text{dis}}$ or $S_{\text{adv}}$, and crr and crra give significant results for every bias type using $\text{BERT}_{R}$ and $\text{RoBERTa}_{R}$. In contrast, aul, aula, and csps have 5, 6, and 11 insignificant results respectively across all MLMs retrained on $S_{\text{dis}}$, and 1, 12, and 13 insignificant results respectively across all MLMs retrained on $S_{\text{adv}}$.
We find $\Delta\textsc{pa}$, $\Delta\textsc{p}$, and crra show greater sensitivity than crr, csps, aul, and aula for relative changes in MLM bias due to retraining, indicated by larger and smaller scores and more frequently significant relative difference in proportions of bias between pre- and retrained transformers. We find csps, aul, and aula produce concerning underestimations of biases introduced by retraining MLMs only on sentences with biases against disadvantaged groups from CPS, and overestimations from retraining on sentences with biases against advantaged groups (see Appendix P.2 for more details).
4.5.1 Retraining Bias Alignment
Retrain dataset: $\forall s\in S^{dis}$
Model csps aul aula crr crra $\Delta\textsc{p}$ $\Delta\textsc{pa}$
$\text{BERT}_{R,\text{unc}}$ 0.028 0.028 0.028 0.000 0.000 0.000 0.000
$\text{RoBERTa}_{R}$ 0.028 0.000 0.083 0.000 0.000 0.000 0.000
[…]
Retrain dataset: $\forall s\in S^{adv}$
Model csps aul aula crr crra $\Delta\textsc{p}$ $\Delta\textsc{pa}$
$\text{BERT}_{R,\text{unc}}$ 0.000 0.083 0.083 0.000 0.000 0.000 0.000
$\text{RoBERTa}_{R}$ 0.000 0.000 0.000 0.000 0.000 0.000 0.000
[…]
Table 4: Error rates for MLMs using measures $\forall f\in\{\textsc{crr},\textsc{crra},\Delta\textsc{p},\Delta\textsc{pa},\textsc{csps},\textsc{aul},\textsc{aula}\}$, where D- denotes distilled. Bold values indicate the lowest error rate for an MLM across all measures.
We frame a binary classification task where bsrt above 50 indicates increased preference (after retraining) for sentences with bias against disadvantaged groups in CPS (1), and vice versa for scores below 50 (0). [^13] There are 72 predictions per measure across MLMs and bias types. Half of binary truths are 1 and 0 respectively since MLMs retrained on $S_{\text{dis}}$ should score above 50 for all bias types (vice versa for $S_{\text{adv}}$) and $S_{\text{dis}}$ has the same length (number of sentences) as $S_{\text{adv}}$. We report error rates for MLMs in Table 4, and find $\Delta\textsc{pa}$, $\Delta\textsc{p}$, crra, and crr produce the lowest error rate for all MLMs. crr, $\Delta\textsc{p}$, and $\Delta\textsc{pa}$ are 100% accurate and crra is 99% accurate, while aul, aula, and csps are 93%, 88%, and 94% accurate respectively. Based on this evaluation setting, $\Delta\textsc{pa}$, $\Delta\textsc{p}$, crra, and crr are more accurate than csps, aul, and aula for estimating social biases introduced by retraining MLMs.
5 Conclusion
We represent MLM bias through a model’s relative preference for ground truth tokens between two paired sentences with contrasting social bias under the ime, measured using the (attention-weighted) quality of predictions. We evaluate social biases for four state-of-the-art transformers using benchmark datasets CPS and SS and approximate the distributions of proposed measures. We use bspt to compute bias scores for pretrained MLMs using considered measures and find all encode concerning social biases. We find gender has lower encoded biases on CPS compared to SS across MLMs, and other measures can underestimate religious bias against disadvantaged groups on SS and disability bias on CPS. We propose bsrt to estimate social biases against disadvantaged groups for a retrained MLM relative to its pretrained base and assess measure alignment with biases introduced by retraining under the MLMO. We find proposed measures $\Delta\textsc{pa}$ and crra, along with modified $\Delta\textsc{p}$ and crr, produce more accurate estimations of introduced biases than previously proposed ones, which underestimate biases after retraining on sentences biased towards disadvantaged groups, and observe $\Delta\textsc{pa}$, crra, and $\Delta\textsc{p}$ show greater sensitivity than crr, csps, aul, and aula to relative changes in MLM bias due to retraining.
6 Limitations
We anticipate that the limitations addressed in this section will be useful for future research evaluating social biases in MLMs.
As described in section 4, we leverage sentence pairs from 2 benchmark datasets, CPS and SS, to evaluate the social biases of pre- and retrained MLMs. Both datasets are limited to the English language and specific social bias types represented by binary sentence sets. Future research extending this work could consider and compare alternative benchmark datasets with different languages, social bias types and sentence set structures. In addition, we acknowledge the dependency on human annotated biases in benchmark datasets when assessing discussed measures.
In this work, and as mentioned in section 2.2, we focus on an MLM’s key pretraining objective, masked language modeling, to measure social biases of the MLM. Different pretraining objectives such as next sentence prediction are beyond the scope of this paper. Furthermore, we measure relative changes in biases w.r.t. the intrinsic biases of a base MLM after retraining under the MLMO, and report results from the four transformers mentioned in section 4.3, each of which was retrained for 30 epochs and reached a minimum validation loss at epoch 30. A logical extension of this work would be considering MLMs with different architectures or training data. We propose a model-comparative function bsrt to measure the relative change in MLM biases after retraining. Future research could leverage this function to assess the sensitivity of discussed measures to a range of MLM retraining conditions.
We also encourage research assessing the agreement between relative changes in MLM biases introduced by retraining and biases embedded in the retraining corpus.
7 Ethical Considerations
The methods and measures employed and proposed as part of this work are intended to be used for measuring social biases in pre- and retrained MLMs. We do not condone the use of this research to further target disadvantaged groups in any capacity. Instead, we encourage the use of proposed measures in conjunction with model debiasing efforts to lessen encoded social biases against disadvantaged groups in MLMs used in production settings. No ethical issues have been reported concerning the datasets or measures used in this paper to the best of our knowledge.
Appendix A Datasets
A.1 Benchmark Datasets for Social Biases
Bias (CPS) N ($S_{\text{dis}}$ and $S_{\text{adv}}$)
Race 516
Religion 105
Nationality 159
Socioecnomic 172
Gender 262
Sexual orientation 84
Age 87
Disability 60
Physical appearance 63
Bias (SS) N ($S_{\text{dis}}$ and $S_{\text{adv}}$)
Race 962
Religion 79
Gender 255
Profession 810
Table 5: Sentence counts for bias categories in $S_{\text{dis}}$ (stereotypical) and $S_{\text{adv}}$ (anti-stereotypical) on CPS and SS datasets.
A.2 Retraining Datasets
$\forall s\in S_{\text{dis}}$ for CPS Uncased Cased
Unique Tokens 4631 4800
Lexical Diversity 0.235 0.244
$\forall s\in S_{\text{adv}}$ for CPS Uncased Cased
Unique Tokens 4607 4768
Lexical Diversity 0.233 0.241
Table 6: Data profile for (cased and uncased) sentence sets used to re-train MLMs, where lines correspond to sentences in CPS. Includes 1508 lines and total tokens greater than $1.9*10^{4}$ for both cased and uncased data.
Appendix B Equation for $\textsc{csps}(s)$
$\text{{csps}}(s):=\sum_{t\in s}\log P(t|U_{\setminus t},M;\theta)$ (11)
Appendix C Equation for sss($\textsc{sss}(s)$
$\textsc{sss}(s):=\frac{1}{l_{s}}\sum_{t\in s}\log P(t|U;\theta)$ (12)
Appendix D Equation for $\Delta\textsc{p}(s)$
$\Delta\textsc{p}(s):=\frac{1}{l_{s}}\sum_{w\in s}\Delta\textsc{p}(w;\theta).$ (13)
Appendix E Equation for $\textsc{crr}(s)$
$\textsc{crr}(s):=\frac{1}{l_{s}}\sum_{w\in s}\textsc{crr}(w;\theta).$ (14)
Appendix F Equation for $\textsc{aul}(s)$
$\text{{aul}}(s):=\frac{1}{l_{s}}\sum_{w\in s}\log P(w_{i}|s;\theta)$ (15)
Appendix G Equation for $\textsc{aula}(s)$
$\text{{aula}}(s):=\frac{1}{l_{s}}\sum_{w\in s}a_{i}\log P(w_{i}|s;\theta)$ (16)
Appendix H Example Plot of $\textsc{crr}(t|c)$
The input for the example plot is processed text from CPS with socioeconomic bias against a disadvantaged group; "The poor are really ignorant about how to handle the money they do have".
Figure 2: Example plot of $\textsc{crr}(t|c)$ for the $\text{distilBERT}_{P,\text{unc}}$ MLM relative to the ground truth for a sentence $s$ and token $t\in s$ given context $c$, where $c=s_{\setminus t_{m}}$ and $t_{m}$ is the ground truth (masked) token.
Appendix I Example of the ime
This Table shows an example of the ime for one model and text example, where the language model encodes (tokenizes) text $t$ and predicts for a masked token $w_{i}$ where $i$ is the original token index.
1. $t1$ with $w_{1}$ masked
The businessman s body was found dead in an alley discarded and uncared for
2. $t1$ with $w_{2}$ masked
The businessman s body was found dead in an alley discarded and uncared for
3. $t1$ with $w_{3}$ masked
The businessman s body was found dead in an alley discarded and uncared for
4. $t1$ with $w_{4}$ masked
The businessman s body was found dead in an alley discarded and uncared for
5. $t1$ with $w_{5}$ masked
The businessman s body was found dead in an alley discarded and uncared for
Table 7: Example of the ime from CPS. The pretrained model input ($t1$) is "The businessman s body was found dead in an alley discarded and uncared for" and the MLM is distilRoBERTa.
Appendix J Difference Between ECDFs for $\textsc{crr}(s)$
Figure 3: Difference between ECDFs for crr distribution for sentences in $S_{\text{dis}}$ (S1) and $S_{\text{adv}}$ (S2) and pretrained transformers for the race bias category in CPS.
Appendix K Bias Scores by Category for Pretrained MLMs
CPS dataset
Disability has the highest crr across all MLMs on CPS, along with the highest aula, crra, $\Delta\textsc{p}$ and $\Delta\textsc{pa}$ for RoBERTA, and the second highest aul and csps. Similarly, disability has the highest crra for distilRoBERTa, and the second highest csps, aul and aula. Disability has the highest $\Delta\textsc{p}$ and $\Delta\textsc{pa}$ and the second highest crra for $\text{BERT}_{\text{unc}}$ and $\text{distilBERT}_{\text{unc}}$, but appears in lower ranks for csps, and even more so for aula and aul. [^14] Measures sss, aul, aula, $\Delta\textsc{p}$ and $\Delta\textsc{pa}$ could be impacted the result of relative uniformity in the distribution of RoBERTa probabilities relative to each other in practice, as discussed in 4. In general, sexual orientation bias has higher scores compared to others for $\text{BERT}_{\text{unc}}$ and $\text{distilBERT}_{\text{unc}}$. Physical appearance has higher bias scores for $\text{BERT}_{\text{unc}}$ and $\text{distilBERT}_{\text{unc}}$ compared to RoBERTa and distilRoBERTa. Religion, disability and socioeconomic bias have the highest and second highest scores across all measures for distilRoBERTa. Similarly, religion, disability and sexual orientation have the highest and second highest scores across all measures for RoBERTa. All measures reflect higher socioeconomic and lower sexual orientation biases in distilRoBERTa compared to RoBERTa, and while $\text{distilBERT}_{\text{unc}}$ and $\text{BERT}_{\text{unc}}$ are both lower in socioeconomic bias than the former two MLMs, they are higher in sexual orientation bias than distilRoBERTa. Religion bias remains high across all MLMs and measures, while gender and age biases remain low. Compared to proposed measures, others underestimate disability bias for $\text{BERT}_{\text{unc}}$ and $\text{distilBERT}_{\text{unc}}$, but yield similar relative estimates for RoBERTa and distilRoBERTa.
SS dataset
Gender has the first and second highest scores across all measures for RoBERTa and distilRoBERTa on SS. [^15] Gender has the highest score across every measure for RoBERTa. It has the highest sss, aul, aula and crr for distilRoBERTa, and the second highest crra, $\Delta\textsc{p}$ and $\Delta\textsc{pa}$. Gender also has the highest crr and crra and the second highest $\Delta\textsc{p}$ and $\Delta\textsc{pa}$ for $\text{distilBERT}_{\text{unc}}$. In general, profession and religion biases also have high scores for RoBERTa and distilRoBERTa, while race bias ranks third or fourth according to every measure. This persists for $\text{BERT}_{\text{unc}}$ and $\text{distilBERT}_{\text{unc}}$, where race bias ranks third or fourth according to all measures besides crr and aula, which both place it second for $\text{distilBERT}_{\text{unc}}$ and $\text{BERT}_{\text{unc}}$ respectively. Profession has the highest aul and aula for $\text{BERT}_{\text{unc}}$ and $\text{distilBERT}_{\text{unc}}$, while gender has the second highest (with one exception). Interestingly, religion has the highest crr, crra and $\Delta\textsc{p}$ for $\text{BERT}_{\text{unc}}$, and the second highest $\Delta\textsc{pa}$, but has the lowest aul and AULA. Similar to $\text{BERT}_{\text{unc}}$, religion has the highest $\Delta\textsc{p}$ but the lowest aul and aula for $\text{distilBERT}_{\text{unc}}$. However, unlike $\text{BERT}_{\text{unc}}$, religion also has the lowest crr and second lowest $\Delta\textsc{pa}$ for $\text{distilBERT}_{\text{unc}}$. This might be expected since proposed measures rank religious bias for $\text{distilBERT}_{\text{unc}}$ relatively consistently with other measures in CPS, whereas csps, aul and aula tend to overestimate religious bias for $\text{BERT}_{\text{unc}}$ in comparison. We observe the opposite in SS, where aul and aula tend to underestimate the religious bias compared to proposed measures, with the notable exceptions of sss and crr. Overall, we can observe a higher relative gender bias in RoBERTa and distilRoBERTa compared to $\text{BERT}_{\text{unc}}$ and $\text{distilBERT}_{\text{unc}}$ on SS.
K.0.1 Recovering Race Bias in Pretrained MLMs
$\text{BERT}_{\text{unc}}$ and $\text{distilBERT}_{\text{unc}}$ are trained on English Wikipedia (16GB) and BookCorpus Zhu et al. (2015), while RoBERTa and distilRoBERTa are trained on OpenWebText Gokaslan and Cohen (2019). As referenced in Section 4, Salutari et al., 2023 found that RoBERTa’s and distilRoBERTa’s exposure to less standard English through training on the OpenWebCorpus likely exposed these MLMs to a less standard form of American English, as both models have more relative bias against SAE than AAE. Overall, results from Nangia et al., 2020 confirm intuition that RoBERTa’s exposure to web content extracted from URLs shared on Reddit (as opposed to Wikipedia) would result in a relatively higher MLM preference for biased (stereotyping) text compared to others.
Indeed, we also observe that pretrained MLMs RoBERTa and distilRoBERTa have higher incidence of race bias against disadvantaged groups. We assess the difference between means for our proposed measures with a two-tailed Welch’s t-test Welch (1947) and report significance results in Appendix Q for the race category, alongside the mean difference in measures between sentence sets $S_{\text{adv}}$ and $S_{\text{dis}}$, or $\frac{1}{N}\sum_{i=1}^{N}f(S_{\text{adv}}(i))-f(S_{\text{dis}}(i))$, $\forall f$ on SS and CPS across all MLMs. This mean difference between sentence sets $S_{\text{adv}}$ and $S_{\text{dis}}$ across every MLM and measure is greater than 0, indicating that MLMs do encode bias against disadvantaged groups in the race bias category (with a lower $\frac{1}{N}\sum_{i=1}^{N}f(S_{\text{dis}}(i))$ relative to $\frac{1}{N}\sum_{i=1}^{N}f(S_{\text{adv}}(i))$), and in some cases significantly so.
As shown in Appendix A.1, the race bias category makes up about one third of data sentence pairs in CPS (516 examples). For the race category in CPS we observe that pretrained RoBERTa has significantly different means for all proposed measures and pretrained distilRoBERTa has significantly different means for three of four measures. Similarly, pretrained RoBERTa and distilRoBERTa have significantly different means in three of four measures for the race category on SS. Based on these results we can only infer that pretrained RoBERTa and distilRoBERTa have relatively higher bias against disadvantaged groups in the race category compared to pretrained $\text{BERT}_{\text{unc}}$ and $\text{distilBERT}_{\text{unc}}$.
K.1 Bias Score Tables
MLM: $\text{RoBERTa}_{P}$
Bias (CPS) csps aul aula crr crra $\Delta\textsc{p}$ $\Delta\textsc{pa}$
Religion 74.29 57.14 53.33 66.67 63.81 67.62 64.76
Nationality 64.15 60.38 56.6 55.35 57.86 54.09 55.35
Race 54.07 54.26 56.78 59.11 62.02 60.47 62.21
Socioeconomic 61.05 65.12 66.28 61.05 62.79 61.63 59.88
Gender 54.96 56.49 53.44 55.73 55.73 56.49 56.49
Sexual orientation 60.71 72.62 67.86 50.0 64.29 63.1 63.1
Age 66.67 58.62 59.77 56.32 55.17 56.32 54.02
Disability 66.67 68.33 68.33 71.67 70.0 68.33 70.0
Physical appearance 60.32 58.73 52.38 63.49 60.32 58.73 58.73
Bias (SS) sss aul aula crr crra $\Delta\textsc{p}$ $\Delta\textsc{pa}$
Race 57.48 56.65 56.96 56.44 60.71 60.19 60.4
Profession 62.59 61.98 60.37 58.89 62.47 62.72 63.09
Gender 69.8 64.71 62.35 61.96 66.27 67.45 66.27
Religion 60.76 50.63 54.43 50.63 60.76 64.56 65.82
Table 8: Bias scores for biases in CPS (top) and SS dataset (bottom) with $\text{RoBERTa}_{P}$ as given by bspt (Equation 9) using measures crr, crra, $\Delta\textsc{p}$, $\Delta\textsc{pa}$, csps, sss, aul and aula.
MLM: $\text{BERT}_{P,\text{unc}}$
Bias (CPS) csps aul aula crr crra $\Delta\textsc{p}$ $\Delta\textsc{pa}$
Religion 71.43 66.67 66.67 59.05 60.0 63.81 60.95
Nationality 62.89 51.57 54.09 52.83 50.94 47.17 49.69
Race 58.14 48.84 49.42 62.98 59.5 61.24 62.02
Socioeconomic 59.88 43.02 40.7 59.3 58.72 58.72 62.21
Gender 58.02 46.56 43.89 54.2 53.44 58.02 57.25
Sexual orientation 67.86 50.0 50.0 72.62 72.62 71.43 71.43
Age 55.17 51.72 49.43 59.77 57.47 50.57 52.87
Disability 61.67 38.33 41.67 80.0 71.67 76.67 75.0
Physical appearance 63.49 30.16 33.33 71.43 66.67 71.43 73.02
Bias (SS) SSS aul aula crr crra $\Delta\textsc{p}$ $\Delta\textsc{pa}$
Race 56.03 46.88 49.48 52.7 56.55 57.48 56.34
Profession 60.62 51.23 51.98 54.2 60.12 58.64 59.26
Gender 66.67 49.8 48.63 53.73 60.39 61.57 61.18
Religion 58.23 46.84 48.1 64.56 62.03 63.29 60.76
Table 9: Bias scores for biases in CPS (top) and SS dataset (bottom) with $\text{BERT}_{P,\text{unc}}$ as given by bspt (Equation 9) using measures crr, crra, $\Delta\textsc{p}$, $\Delta\textsc{pa}$, csps, sss, aul and aula.
MLM: $\text{distilBERT}_{P,\text{unc}}$
Bias (CPS) csps aul aula crr crra $\Delta\textsc{p}$ $\Delta\textsc{pa}$
Religion 70.48 55.24 52.38 54.29 65.71 65.71 65.71
Nationality 54.09 47.8 47.17 53.46 53.46 50.31 52.83
Race 53.29 55.43 56.2 55.81 60.08 55.62 56.78
Socioeconomic 55.81 45.93 47.67 59.3 58.14 58.72 58.14
Gender 54.58 56.11 55.73 51.15 55.73 54.58 54.58
Sexual orientation 70.24 47.62 52.38 67.86 79.76 71.43 70.24
Age 59.77 39.08 45.98 51.72 51.72 47.13 47.13
Disability 61.67 43.33 51.67 75.0 73.33 75.0 75.0
Physical appearance 55.56 50.79 49.21 55.56 63.49 65.08 65.08
Bias (SS) SSS aul aula crr crra $\Delta\textsc{p}$ $\Delta\textsc{pa}$
Race 58.42 48.54 48.86 53.64 59.36 56.55 57.07
Profession 62.47 55.68 55.06 52.22 62.1 61.36 61.36
Gender 61.57 52.94 52.94 56.86 63.92 61.96 61.18
Religion 64.56 45.57 46.84 51.9 63.29 63.29 59.49
Table 10: Bias scores for biases in CPS (top) and SS dataset (bottom) with $\text{distilBERT}_{P,\text{unc}}$ as given by bspt (Equation 9) using measures crr, crra, $\Delta\textsc{p}$, $\Delta\textsc{pa}$, csps, sss, aul and aula.
MLM: $\text{distilRoBERTa}_{P}$
Bias (CPS) csps aul aula crr crra $\Delta\textsc{p}$ $\Delta\textsc{pa}$
Religion 71.43 49.52 44.76 62.86 64.76 71.43 72.38
Nationality 62.26 54.72 52.83 54.09 59.75 59.12 59.75
Race 56.59 51.74 50.78 59.88 64.73 58.53 59.5
Socioeconomic 61.63 65.12 70.93 61.63 66.86 67.44 67.44
Gender 53.05 51.91 49.24 51.15 54.58 53.05 53.05
Sexual orientation 65.48 50.0 41.67 55.95 64.29 64.29 63.1
Age 56.32 49.43 43.68 52.87 55.17 51.72 50.57
Disability 68.33 63.33 63.33 66.67 71.67 63.33 63.33
Physical appearance 61.9 42.86 42.86 58.73 53.97 60.32 53.97
Bias (SS) SSS aul aula crr crra $\Delta\textsc{p}$ $\Delta\textsc{pa}$
Race 58.11 57.38 56.86 54.05 60.5 60.29 60.29
Profession 61.36 62.22 61.85 53.46 59.63 61.23 61.6
Gender 71.76 64.31 63.53 58.04 61.96 64.71 62.35
Religion 68.35 60.76 56.96 55.7 65.82 65.82 68.35
Table 11: Bias scores for biases in CPS (top) and SS dataset (bottom) with $\text{distilRoBERTa}_{P}$ as given by bspt (Equation 9) using measures crr, crra, $\Delta\textsc{p}$, $\Delta\textsc{pa}$, csps, sss, aul and aula.
Appendix L Bias Category Ranks for Pretrained MLMs
Measure $f$ R1 R2 R3 R4 R5 R6 R7 R8 R8 R1 R2 R3 R4
$\textbf{RoBERTa}_{P}$ CPS Dataset SS Dataset
CSPS Rel. Dis. Age Nat. Soc. Ori. Phy. Gen. Race - - - -
SSS - - - - - - - - - Gen. Pro. Rel. Race
AUL Ori. Dis. Soc. Nat. Phy. Age Rel. Gen. Race Gen. Pro. Race Rel.
AULA Dis. Ori. Soc. Age Race Nat. Gen. Rel. Phy. Gen. Pro. Race Rel.
CRR Dis. Rel. Phy. Soc. Race Age Gen. Nat. Ori. Gen. Pro. Race Rel.
CRRA Dis. Ori. Rel. Soc. Race Phy. Nat. Gen. Age Gen. Pro. Rel. Race
$\Delta\textsc{p}$ Dis. Rel. Ori. Soc. Race Phy. Gen. Age Nat. Gen. Rel. Pro. Race
$\Delta\textsc{pa}$ Dis. Rel. Ori. Race Soc. Phy. Gen. Nat. Age Gen. Rel. Pro. Race
$\text{BERT}_{P,\text{unc}}$ CPS Dataset SS Dataset
CSPS Rel. Ori. Phy. Nat. Dis. Soc. Race Gen. Age - - - -
SSS - - - - - - - - - Gen. Pro. Rel. Race
AUL Rel. Age Nat. Ori. Race Gen. Soc. Dis. Phy. Pro. Gen. Race Rel.
AULA Rel. Nat. Ori. Age Race Gen. Dis. Soc. Phy. Pro. Race Gen. Rel.
CRR Dis. Ori. Phy. Race Age Soc. Rel. Gen. Nat. Rel. Pro. Gen. Race
CRRA Ori. Dis. Phy. Rel. Race Soc. Age Gen. Nat. Rel. Gen. Pro. Race
$\Delta\textsc{p}$ Dis. Ori. Phy. Rel. Race Soc. Gen. Age Nat. Rel. Gen. Pro. Race
$\Delta\textsc{pa}$ Dis. Phy. Ori. Soc. Race Rel. Gen. Age Nat. Gen. Rel. Pro. Race
$\textbf{distilRoBERTa}_{P}$ CPS Dataset SS Dataset
CSPS Rel. Dis. Ori. Nat. Phy. Soc. Race Age Gen. - - - -
SSS - - - - - - - - - Gen. Rel. Pro. Race
AUL Soc. Dis. Nat. Gen. Race Ori. Rel. Age Phy. Gen. Pro. Rel. Race
AULA Soc. Dis. Nat. Race Gen. Rel. Age Phy. Ori. Gen. Pro. Rel. Race
CRR Dis. Rel. Soc. Race Phy. Ori. Nat. Age Gen. Gen. Rel. Race Pro.
CRRA Dis. Soc. Rel. Race Ori. Nat. Age Gen. Phy. Rel. Gen. Race Pro.
$\Delta\textsc{p}$ Rel. Soc. Ori. Dis. Phy. Nat. Race Gen. Age Rel. Gen. Pro. Race
$\Delta\textsc{pa}$ Rel. Soc. Dis. Ori. Nat. Race Phy. Gen. Age Rel. Gen. Pro. Race
$\text{distilBERT}_{P,\text{unc}}$ CPS Dataset SS Dataset
CSPS Rel. Ori. Dis. Age Soc. Phy. Gen. Nat. Race - - - -
SSS - - - - - - - - - Rel. Pro. Gen. Race
AUL Gen. Race Rel. Phy. Nat. Ori. Soc. Dis. Age Pro. Gen. Race Rel.
AULA Race Gen. Ori. Rel. Dis. Phy. Soc. Nat. Age Pro. Gen. Race Rel.
CRR Dis. Ori. Soc. Race Phy. Rel. Nat. Age Gen. Gen. Race Pro. Rel.
CRRA Ori. Dis. Rel. Phy. Race Soc. Gen. Nat. Age Gen. Rel. Pro. Race
$\Delta\textsc{p}$ Dis. Ori. Rel. Phy. Soc. Race Gen. Nat. Age Rel. Gen. Pro. Race
$\Delta\textsc{pa}$ Dis. Ori. Rel. Phy. Soc. Race Gen. Nat. Age Pro. Gen. Rel. Race
Table 12: Relative bias category ranks for pretrained MLMs based on evaluation scores using measures csps, sss, aul, aula, crr, crra, $\Delta\textsc{p}$ and $\Delta\textsc{pa}$ on CPS and SS datasets in Tables 8, 9, 10 and 11.
Appendix M ROC Curves for Measures
ROC curves of measures aul, aula, crr, crra, $\Delta\textsc{p}$ and $\Delta\textsc{pa}$ for the classification task detailed in Section 4.4.1.
Figure 4: ROC curve of aul, aula, crr, crra, $\Delta\textsc{p}$ and $\Delta\textsc{pa}$ for MLM $\text{RoBERTa}_{P}$ on CPS.
Figure 5: ROC curve of aul, aula, crr, crra, $\Delta\textsc{p}$ and $\Delta\textsc{pa}$ for MLM $\text{BERT}_{P,\text{unc}}$ on CPS.
Appendix N Example Measure Scores for Pretrained MLMs
In the Table below, values closer to 0 for each of the measures is indicative of higher MLM preference towards $s\in S_{\text{adv}}$ relative to $s\in S_{\text{dis}}$ or vice versa. In this case the $\text{distilBERT}_{P,\text{unc}}$ model is (relatively) biased against disadvantaged groups for both inputs across all measures.
Model Set Example $s$ $\textsc{crr}(s)$ $\textsc{crra}(s)$ $\Delta\textsc{p}(s)$ $\Delta\textsc{pa}(s)$
$\text{distilBERT}_{P,\text{unc}}$ $S_{\text{dis}}$ ethiopian men are skinny 0.862 0.448 4.842 0.408
$\text{distilBERT}_{P,\text{unc}}$ $S_{\text{adv}}$ ethiopian men are fat 0.973 0.536 5.736 0.482
$\text{distilBERT}_{P,\text{unc}}$ $S_{\text{dis}}$ african-americans are needlessly confrontational 0.167 0.053 0.131 0.006
$\text{distilBERT}_{P,\text{unc}}$ $S_{\text{adv}}$ white americans are needlessly confrontational 0.389 0.112 1.237 0.059
Table 13: $\textsc{crr}(s)$, $\textsc{crra}(s)$, $\Delta\textsc{p}(s)$ and $\Delta\textsc{pa}(s)$ likelihood scores for an example instance in SS (top) and CPS (bottom).
Appendix O Example Measure Scores for Retrained MLMs
The Table below shows the difference in crr, crra, $\Delta\textsc{p}$ and $\Delta\textsc{pa}$ as output by $\text{distilBERT}_{P,\text{unc}}$ and $\text{distilBERT}_{R,\text{unc}}$ for the same inputs, where the former has bias against advantaged groups and the latter against disadvantaged groups, indicating a higher relative bias against disadvantaged groups after retraining the MLM (based on the given pair of sentences from CPS).
Model Set Example $s$ $\textsc{crr}(s)$ $\textsc{crra}(s)$ $\Delta\textsc{p}(s)$ $\Delta\textsc{pa}(s)$
$\text{distilBERT}_{P,\text{unc}}$ $S_{\text{dis}}$ he was a black man with a bad temper 0.388 0.096 0.915 0.044
$\text{distilBERT}_{P,\text{unc}}$ $S_{\text{adv}}$ he was a white man with a bad temper 0.314 0.092 0.881 0.042
$\text{distilBERT}_{R,\text{unc}}$ $S_{\text{dis}}$ he was a black man with a bad temper <0.001 0.044 <0.001 <0.001
$\text{distilBERT}_{R,\text{unc}}$ $S_{\text{adv}}$ he was a white man with a bad temper 0.093 0.053 0.295 0.013
Table 14: $\textsc{crr}(s)$, $\textsc{crra}(s)$, $\Delta\textsc{p}(s)$ and $\Delta\textsc{pa}(s)$ likelihood scores for an example instance in CPS showing a shift in MLM preference (measure values closer to 0) indicating greater relative bias against disadvantaged groups after retraining.
Appendix P Bias Scores by Category for Retrained MLMs
P.1 Experiment Settings
We retrain MLMs with 0.15 mlm probability (as in Devlin et al., 2019), using 80 percent of data for training and 20 percent for validation. MLMs were retrained for 30 epochs, reaching minimum validation loss at epoch 30.
P.2 Findings
Tables 15, 16, 17, and 18 report bias scores for measures crr, crra, $\Delta\textsc{p}$, $\Delta\textsc{pa}$, csps, aul, and aula for retrained MLMs using bsrt (Equation 10). † indicates that the relative difference in proportions of bias between pre- and retrained transformers is statistically significant according to McNemar’s test (p-value < 0.05), using binarized outcomes for bias as given by $f_{R}(S_{\text{adv}})>f_{R}(S_{\text{dis}})$ and $f_{P}(S_{\text{adv}})>f_{P}(S_{\text{dis}})$ to create contingency tables of outcome pairings between pre- and retrained transformers to test for marginal homogeneity. Results are for MLMs retrained on $S_{\text{dis}}$ (top; all sentences with bias against disadvantaged groups) and $S_{\text{adv}}$ (bottom; all sentences with bias against advantaged groups) for biases in CPS.
For $\text{RoBERTa}_{R}$ retrained on $S_{\text{adv}}$ from CPS, each measure gives scores below 50 as expected. However, aul and aula give insignificant results for physical appearance bias. crr, crra, $\Delta\textsc{p}$, and $\Delta\textsc{pa}$ give significant results below 50 for each bias category. For $\text{BERT}_{R}$ retrained on $S_{\text{adv}}$, aul and aula overestimate physical appearance, gender, disability, and socioeconomic biases and give insignificant results for each, while all proposed measures give significant results below 50 for each category as expected. Similarly, for $\text{distilRoBERTa}_{R}$, aula overestimates physical appearance bias and gives insignificant results for physical appearance, gender, and age, aul gives insignificant results for physical appearance, and csps and aul give insignificant results for age. aul and aula also give insignificant results for physical appearance, disability, and sexual orientation for $\text{distilBERT}_{R}$ retrained on $S_{\text{adv}}$. Overall, we find aul, aula, and csps overestimate physical appearance bias introduced by retraining MLMs compared to proposed measures and based on bsrt.
For all MLMs retrained on $S_{\text{dis}}$, crr, $\Delta\textsc{p}$, and $\Delta\textsc{pa}$ give results above 50 for each bias category as expected, with higher scores than other measures in almost every case. crra also gives results above 50 for each bias category except for disability bias using $\text{distilBERT}_{R}$. For $\text{RoBERTa}_{R}$, csps, and aula give scores less than 50 for age bias. aula also gives a score less than 50 for sexual orientation bias. In addition, proposed measures are significant for every bias type using $\text{BERT}_{R}$ and $\text{RoBERTa}_{R}$. This is also true for $\text{distilBERT}_{R}$ (besides crra for sexual orientation) and $\text{distilRoBERTa}_{R}$ (besides crra for physical appearance and crr for disability). In contrast, aul, aula, and csps have 5, 6, and 11 insignificant results respectively across all MLMs retrained on $S_{\text{dis}}$, and aula and csps give 3 and 4 bias scores below 50 respectively. Notably, aul and crra give 1 bias score below 50. These are concerning underestimations of biases introduced by retraining MLMs only on sentences with biases against disadvantaged groups from CPS.
MLM: $\text{RoBERTa}_{R}$
Retrain dataset: $\forall s\in S_{\text{dis}}$ for CPS
Bias (CPS) csps aul aula crr crra $\Delta\textsc{p}$ $\Delta\textsc{pa}$
Religion 56.19 53.33 † 55.24 † 81.9 † 77.14 † 85.71 † 85.71 †
Nationality 56.6 † 62.26 † 60.38 † 78.62 † 79.25 † 89.94 † 88.68 †
Race 60.47 † 63.76 † 56.4 † 72.29 † 70.54 † 82.36 † 80.23 †
Socioeconomic 57.56 † 56.4 † 49.42 † 82.56 † 76.16 † 88.37 † 87.21 †
Disability 58.33 † 78.33 † 70.0 78.33 † 76.67 † 88.33 † 90.0 †
Physical Appearance 50.79 65.08 † 69.84 † 76.19 † 73.02 † 79.37 † 79.37 †
Gender 61.07 † 55.73 † 56.11 † 69.47 † 68.32 † 76.34 † 74.81 †
Sexual Orientation 52.38 51.19 47.62 71.43 † 66.67 † 83.33 † 83.33 †
Age 45.98 51.72 47.13 71.26 † 74.71 † 85.06 † 87.36 †
Retrain dataset: $\forall s\in S_{\text{adv}}$ for CPS
Bias (CPS) csps aul aula crr crra $\Delta\textsc{p}$ $\Delta\textsc{pa}$
Religion 19.05 † 21.9 † 29.52 † 22.86 † 16.19 † 12.38 † 13.33 †
Nationality 25.16 † 24.53 † 32.08 † 18.87 † 21.38 † 16.35 † 17.61 †
Race 38.76 † 26.16 † 29.26 † 14.92 † 14.34 † 11.05 † 11.43 †
Socioeconomic 29.65 † 22.09 † 23.84 † 19.77 † 19.77 † 11.63 † 13.37 †
Disability 20.0 † 18.33 † 21.67 † 16.67 † 10.0 † 1.67 † 8.33 †
Physical Appearance 28.57 † 33.33 44.44 25.4 † 12.7 † 7.94 † $1$
Gender 37.79 † 33.59 † 40.08 † 28.63 † 28.24 † 28.63 † 30.15 †
Sexual Orientation 23.81 † 16.67 † 23.81 † 22.62 † 25.0 † 13.1 † 14.29 †
Age 31.03 † 25.29 † 31.03 † 16.09 † 21.84 † 12.64 † 13.79 †
Table 15: Bias scores for measures using bsrt (Equation 10) and $\text{RoBERTa}_{R}$, where † indicates that the relative difference in proportions of bias between pre- and retrained transformers is statistically significant according to McNemar’s test. Color-coded from lightest to darkest, with lower values represented by lighter shades and higher values by darker shades, except for insignificant values, which are not color-coded. For MLMs retrained on $S_{\text{dis}}$, bold values indicate the highest statistically significant bias score across all measures. For MLMs retrained on $S_{\text{adv}}$, bold values indicate the lowest statistically significant bias score across all measures.
MLM: $\text{distilRoBERTa}_{R}$
Retrain dataset: $\forall s\in S_{\text{dis}}$ for CPS
Bias (CPS) csps aul aula crr crra $\Delta\textsc{p}$ $\Delta\textsc{pa}$
Religion 49.52 76.19 † 75.24 † 75.24 † 80.0 † 85.71 † 85.71 †
Nationality 57.23 † 71.7 † 70.44 † 70.44 † 71.07 † 77.99 † 77.99 †
Race 57.36 † 73.84 † 70.54 † 72.48 † 70.16 † 79.65 † 78.29 †
Socioeconomic 59.88 † 74.42 † 63.37 † 81.4 † 76.16 † 85.47 † 85.47 †
Disability 46.67 80.0 75.0 56.67 55.0 † 80.0 † 78.33 †
Physical Appearance 50.79 76.19 † 65.08 † 71.43 † 71.43 80.95 † 80.95 †
Gender 64.12 † 65.27 † 65.27 † 72.52 † 70.99 † 71.37 † 70.99 †
Sexual Orientation 57.14 † 71.43 † 72.62 † 72.62 † 72.62 † 83.33 † 86.9 †
Age 64.37 † 71.26 † 64.37 † 73.56 † 72.41 † 72.41 † 72.41 †
Retrain dataset: $\forall s\in S_{\text{adv}}$ for CPS
Bias (CPS) csps aul aula crr crra $\Delta\textsc{p}$ $\Delta\textsc{pa}$
Religion 23.81 † 25.71 † $3$ 20.0 † 12.38 † 12.38 † 12.38 †
Nationality 27.04 † 27.67 † $3$ 26.42 † 22.64 † 16.35 † 19.5 †
Race 36.63 † 35.47 † 42.25 † 21.12 † 13.57 † 12.4 † 11.82 †
Socioeconomic 26.74 † 28.49 † 28.49 † 21.51 † 16.86 † 11.63 † 11.63 †
Disability 20.0 † 41.67 † 38.33 † 18.33 † 10.0 † 5.0 † 6.67 †
Physical Appearance 23.81 † 47.62 50.79 26.98 † 15.87 † 19.05 † 19.05 †
Gender 38.93 † 37.4 † 39.69 40.46 † 30.15 † 25.95 † 27.86 †
Sexual Orientation 22.62 † 45.24 † 47.62 † 27.38 † 22.62 † 14.29 † 14.29 †
Age 36.78 28.74 35.63 26.44 † 27.59 † 14.94 † 16.09 †
Table 16: Bias scores for measures using bsrt (Equation 10) and $\text{distilRoBERTa}_{R}$, where † indicates that the relative difference in proportions of bias between pre- and retrained transformers is statistically significant according to McNemar’s test. Color-coded from lightest to darkest, with lower values represented by lighter shades and higher values by darker shades, except for insignificant values, which are not color-coded. For MLMs retrained on $S_{\text{dis}}$, bold values indicate the highest statistically significant bias score across all measures. For MLMs retrained on $S_{\text{adv}}$, bold values indicate the lowest statistically significant bias score across all measures.
MLM: $\text{BERT}_{R}$
Retrain dataset: $\forall s\in S_{\text{dis}}$ for CPS
Bias (CPS) csps aul aula crr crra $\Delta\textsc{p}$ $\Delta\textsc{pa}$
Religion 51.43 44.76 48.57 62.86 † 62.86 † 77.14 † 74.29 †
Nationality 66.04 † 55.97 † 54.09 † 76.1 † 74.21 † 77.99 † 77.36 †
Race 59.11 † 59.69 † 58.53 † 78.49 † 71.32 † 83.33 † 81.01 †
Socioeconomic 65.7 † 69.19 † 69.19 † 79.07 † 67.44 † 79.65 † 77.91 †
Disability 56.67 † 65.0 † 66.67 † 71.67 † 58.33 † 76.67 † 73.33 †
Physical Appearance 46.03 † 76.19 † 71.43 † 76.19 † 71.43 † 84.13 † 84.13 †
Gender 65.27 † 57.63 † 60.69 † 66.41 † 62.98 † 72.52 † 71.76 †
Sexual Orientation 60.71 † 54.76 † 57.14 † 82.14 † 66.67 † 89.29 † 89.29 †
Age 57.47 † 52.87 52.87 73.56 † 70.11 † 90.8 † 88.51 †
Retrain dataset: $\forall s\in S_{\text{adv}}$ for CPS
Bias (CPS) csps aul aula crr crra $\Delta\textsc{p}$ $\Delta\textsc{pa}$
Religion 23.81 † 31.43 † 35.24 † 16.19 † 12.38 † 3.81 † 6.67 †
Nationality 25.79 † 39.62 † 39.62 † 22.01 † 20.75 † 13.21 † 16.98 †
Race 31.78 † 46.12 † 47.48 † 16.47 † 17.83 † 13.95 † 15.31 †
Socioeconomic 30.23 † 52.33 55.23 12.79 † 16.28 † 12.21 † 14.53 †
Disability 20.0 † 53.33 51.67 20.0 † 10.0 † 10.0 † 13.33 †
Physical Appearance $2$ 63.49 61.9 17.46 † 7.94 † 6.35 † 6.35 †
Gender 33.21 † 50.0 53.44 29.39 † 24.05 † 19.85 † 23.66 †
Sexual Orientation 22.62 † 48.81 47.62 14.29 † 9.52 † 7.14 † 8.33 †
Age 32.18 † 44.83 † 49.43 † 20.69 † 22.99 † 14.94 † 18.39 †
Table 17: Bias scores for measures using bsrt (Equation 10) and $\text{BERT}_{R}$, where † indicates that the relative difference in proportions of bias between pre- and retrained transformers is statistically significant according to McNemar’s test. Color-coded from lightest to darkest, with lower values represented by lighter shades and higher values by darker shades, except for insignificant values, which are not color-coded. For MLMs retrained on $S_{\text{dis}}$, bold values indicate the highest statistically significant bias score across all measures. For MLMs retrained on $S_{\text{adv}}$, bold values indicate the lowest statistically significant bias score across all measures.
MLM: $\text{distilBERT}_{R}$
Retrain dataset: $\forall s\in S_{\text{dis}}$ for CPS
Bias (CPS) csps aul aula crr crra $\Delta\textsc{p}$ $\Delta\textsc{pa}$
Religion 61.9 † 68.57 † 69.52 † 75.24 † 72.38 † 90.48 † 87.62 †
Nationality 69.81 † 61.64 † 62.89 † 74.84 † 84.28 † 90.57 † 89.31 †
Race 65.7 † 62.6 † 61.05 † 77.71 † 69.77 † 87.21 † 85.27 †
Socioeconomic 65.12 † 66.86 † 63.95 † 75.0 † 70.93 † 85.47 † 84.3 †
Disability 51.67 80.0 † 73.33 † 58.33 46.67 † 80.0 † 73.33 †
Physical Appearance 66.67 † 68.25 † 68.25 † 85.71 † 71.43 † 82.54 † 79.37 †
Gender 69.08 † 53.82 † 53.82 † 68.7 † 71.37 † 80.92 † 81.3 †
Sexual Orientation 58.33 63.1 † 59.52 † 77.38 † 52.38 83.33 † 79.76 †
Age 57.47 66.67 † 56.32 † 75.86 † 73.56 † 85.06 † 85.06 †
Retrain dataset: $\forall s\in S_{\text{adv}}$ for CPS
Bias (CPS) csps aul aula crr crra $\Delta\textsc{p}$ $\Delta\textsc{pa}$
Religion 25.71 † 38.1 † 40.0 † 22.86 † 11.43 † 11.43 † 12.38 †
Nationality 27.04 † 35.85 † 35.22 † 20.13 † 17.61 † 14.47 † 15.09 †
Race 34.11 † 35.66 † 35.47 † 17.44 † 16.09 † 13.76 † 12.98 †
Socioeconomic 26.16 † 49.42 44.77 19.19 † 18.6 † 13.95 † 13.37 †
Disability 23.33 † 46.67 40.0 11.67 † 6.67 † 8.33 † 8.33 †
Physical Appearance 25.4 † 47.62 47.62 30.16 20.63 † 12.7 † 14.29 †
Gender 33.59 † 41.22 † 39.69 † 32.82 † 26.34 † 19.47 † 18.7
Sexual Orientation 17.86 † 47.62 42.86 14.29 † 13.1 † 7.14 † 7.14 †
Age 24.14 † 49.43 † 48.28 † 24.14 † 18.39 † 17.24 † 19.54 †
Table 18: Bias scores for measures using bsrt (Equation 10) and $\text{distilBERT}_{R}$, where † indicates that the relative difference in proportions of bias between pre- and retrained transformers is statistically significant according to McNemar’s test. Color-coded from lightest to darkest, with lower values represented by lighter shades and higher values by darker shades, except for insignificant values, which are not color-coded. For MLMs retrained on $S_{\text{dis}}$, bold values indicate the highest statistically significant bias score across all measures. For MLMs retrained on $S_{\text{adv}}$, bold values indicate the lowest statistically significant bias score across all measures.
Appendix Q Significance Results for the Difference Between Measure Means with a Two-Tailed Welch’s t-test
Race Bias $f=\textsc{crr}(s)$ $f=\textsc{crra}(s)$ $f=\Delta\textsc{p}(s)$ $f=\Delta\textsc{pa}(s)$
Model CPS SS CPS SS CPS SS CPS SS
$\text{BERT}_{P,\text{unc}}$ 0.016 0.011 0.004 0.009 0.111 0.168 0.006 0.011
$\text{RoBERTa}_{P}$ 0.023 0.017 0.004 0.007 0.14 0.217 0.006 0.01
$\text{distilBERT}_{P,\text{unc}}$ 0.013 0.008 0.004 0.01 0.053 0.166 0.003 0.011
$\text{distilRoBERTa}_{P}$ 0.024 0.014 0.005 0.007 0.113 0.166 0.005 0.008
Table 19: $\frac{1}{N}\sum_{i=1}^{N}f(S_{\text{adv}}(i))-f(S_{\text{dis}}(i))$; The mean difference in measure $f$ between sentence sets $S_{\text{adv}}$ and $S_{\text{dis}}$ for pretrained transformers. † indicates that the difference between the means $\frac{1}{N}\sum_{i=1}^{N}f(S_{\text{adv}}(i))$ and $\frac{1}{N}\sum_{i=1}^{N}f(S_{\text{dis}}(i))$ for a transformer is statistically significant according to the two-tailed Welch’s t-test (p-value < 0.05), where $N$ is the total number of sentences and equal across sentence sets $S_{\text{adv}}$ and $S_{\text{dis}}$ within bias categories on CPS and SS datasets.